# Dirac particles, spin and photons

### v1

@@ 1. Introduction
In fact, already at the level of classical mechanics one should understand what the spin of a particle is and what an antiparticle is. The situation with spin is as follows. The relativistic phase space $T^{*}\mathbb{R}^{1,3}$ of a massive particle with fixed parity extends to a phase space $T^{*}\mathbb{R}^{1,3}\times\mathbb{C}^{2}$, where $\mathbb{C}^{2}$ is the fundamental representation of the group $\mathrm{SL}(2,\mathbb{C})$ doubly covering the proper ortochronous Lorentz group $\mathrm{SO}^{+}(1,3)$ [1]. Next, a formal transition is made to the orbit $T^{*}H_{+}^{3}\times{\mathbb{C}P}^{1}$ of the Poincare group, over this orbit a complex line bundle $L_{\mathbb{C}}^{+}$ is introduced and the polarized sections of this bundle define fields with spin $s=0,{\textstyle\frac{1}{2}},1,$... [1]. Here $H_{+}^{3}$ is one sheet of a two-sheeted hyperboloid $\eta^{\mu\nu}p_{\mu}p_{\nu}-m^{2}=0$ in the momentum space, $p_{\mu}\in\mathbb{R}^{1,3}$, $\mu,\nu,...=0,...,3$, $(\eta_{\mu\nu})={\mathrm{diag}}(1,-1,-1,-1)$. If the parity is not fixed, then the phase space should be extended to the space $T^{*}\mathbb{R}^{1,3}\times\mathbb{C}^{2}_{L}\times\mathbb{C}_{R}^{2}$, where $\mathbb{C}_{L}^{2}$ and $\mathbb{C}_{R}^{2}$ are the representation spaces of the Lorentz group of type $({\textstyle\frac{1}{2}},0)$ and $(0,{\textstyle\frac{1}{2}})$. This allows us to obtain a description of fields in arbitrary representations $(s,j)$ of the Lorentz group, and not only representations of type $(s,0)$ and $(0,s)$ obtained for fixed parity. In the non-relativistic case, the phase space $\mathbb{C}^{2}_{L}\times\mathbb{C}_{R}^{2}$ is reduced to the fundamental representation $\mathbb{C}^{2}$ of the group SU(2) covering the rotation group SO(3), and the coset space ${\mathbb{C}P}^{1}=\mathrm{SU}(2)/\mathrm{U}(1)\subset\mathbb{C}^{2}$ of this group is used to introduce the space $\mathbb{C}^{2s+1}$ of quantum spin. The space of quantum spin in both the relativistic and non-relativistic cases is defined as the space $\mathbb{C}^{2s+1}$ of sections of the bundle $\mathcal{O}(2s)$ over the space ${\mathbb{C}P}^{1}$.
@@ 2.1. Dynamical spin variables
$N_{int}^{L}=\frac{p^{{{\alpha}}\dot{{\alpha}}}}{m}z_{{\alpha}}\bar{z}_{\dot{{\alpha}}}=2s\ ,$ (2.6)
@@ 2.3. Spinning particles
$V_{S^{a}}=\mathrm{i}\sigma^{{\alpha}}_{a\beta}\left(z_{{\alpha}}\frac{\partial}{\partial z_{\beta}}+\hat{z}^{{\alpha}}\frac{\partial}{\partial\hat{z}_{\beta}}\right)\quad\Rightarrow\quad[V_{S^{a}},V_{S^{b}}]={\varepsilon}^{c}_{ab}V_{S^{c}}\ ,$ (2.37)
@@ 2.5. Higher spins
$\mathrm{d}s^{2}_{\mathbb{C}^{2}/\mathbb{Z}_{n}}=\mathrm{d}\rho^{2}+\rho^{2}\mathrm{d}s^{2}_{S^{3}/\mathbb{Z}_{n}}=\mathrm{d}\rho^{2}+\frac{\rho^{2}}{n^{2}}(\mathrm{d}{{\varphi}}-\mathrm{i}a_{-n})^{2}+\frac{\mathrm{d}z\wedge\mathrm{d}{\bar{z}}}{(1+z{\bar{z}})^{2}}\ .$ (2.53)
@@
$\mathrm{d}s^{2}_{\mathbb{C}/\mathbb{Z}_{m}}=\mathrm{d}z\mathrm{d}{\bar{z}}|_{\mathbb{C}/\mathbb{Z}_{m}}=\frac{1}{m^{2}}(\psi_{m}^{*}\psi_{m})^{\frac{1-m}{m}}\mathrm{d}\psi_{n}\mathrm{d}\psi_{m}^{*}=\mathrm{d}\rho^{2}+\frac{\rho^{2}}{m^{2}}\mathrm{d}{{\varphi}}^{2}\ ,$ (2.56)
@@ 3.3. Schrödinger equations
$\nabla_{x^{a}}=\frac{\partial}{\partial x^{a}}+p_{a}{J_{\sf{v}}}\ ,\quad\nabla_{p_{a}}=\frac{\partial}{\partial p_{a}}\quad\Rightarrow\quad\hat{x}^{a}={J_{\sf{v}}}\nabla_{p_{a}}{\quad\mathrm{and}\quad}\hat{p}_{a}=-\mathrm{i}\nabla_{x^{a}}\ .$ (3.24)
@@
$\nabla_{x^{a}}^{\pm}=\pm\mathrm{i}p_{a}\ ,\quad\nabla_{p_{a}}^{\pm}=\frac{\partial}{\partial p_{a}}\quad\Rightarrow\quad\hat{x}^{a}_{\pm}=\pm\mathrm{i}\frac{\partial}{\partial p_{a}}{\quad\mathrm{and}\quad}\hat{p}_{a}^{\pm}=\pm p_{a}\ ,$ (3.25)
@@ 3.4. Wave function collapse
Thus, the probability of detecting certain parameters $(s,s_{3})$ in an experiment arises in the case where we do not know fully the initial conditions of the particle’s motion. This is equivalent to how probabilities arise when throwing dice.
@@ 3.5. From Schrödinger to Pauli equation
$\langle\Psi_{\pm},\Psi\pm\rangle=\Phi_{\pm}^{\dagger}\Phi_{\pm}=\delta_{{{\alpha}}{{\dot{\alpha}}}}\psi_{\pm}^{{\alpha}}\bar{\psi}_{\pm}^{{\dot{\alpha}}}{\quad\mathrm{and}\quad}\rho_{\pm}=\pm\Phi_{\pm}^{\dagger}\Phi_{\pm}$ (3.57)
@@ 4. Relativistic Hamiltonian mechanics
4.1. Free spinless particles
@@ 4.4. Dirac and Proca particles
$g_{D}=-\mathrm{i}\,\mathrm{d}Z^{\pm}J\mathrm{d}\bar{Z}^{\pm}=\mathrm{d}Z^{+}\wedge\mathrm{d}\bar{Z}^{+}=-\mathrm{d}Z^{-}\wedge\mathrm{d}\bar{Z}^{-}=\mathrm{d}z_{{\alpha}}\mathrm{d}\bar{y}^{{\alpha}}+\mathrm{d}y^{{\dot{\alpha}}}\mathrm{d}{\bar{z}}_{{\dot{\alpha}}}\ ,$ (4.60)
@@ 5.2. Relativistic analogue of Schrödinger equation
$\begin{split}L_{\mathbb{C}}^{\pm}&\ \to\ T^{*}\mathbb{R}^{1,3}:\quad\nabla_{x^{\mu}}=\partial_{\mu}\pm\mathrm{i}p_{\mu}\ ,\quad\nabla_{p_{\mu}}=\frac{\partial}{\partial p_{\mu}}\ ,\\[2.0pt] L_{\mathbb{C}}^{\pm}&\ \to\ \mathbb{C}_{\pm}^{4}:\quad\begin{array}[]{l}\nabla_{z_{{\alpha}}^{\pm}}=\partial_{z_{{\alpha}}^{\pm}}+{\textstyle\frac{1}{2}}v^{{{\alpha}}{{\dot{\alpha}}}}{\bar{z}}_{{\dot{\alpha}}}^{\pm}\ ,\ \nabla_{{\bar{z}}_{{\dot{\alpha}}}^{\pm}}=\partial_{{\bar{z}}_{{\dot{\alpha}}}^{\pm}}-{\textstyle\frac{1}{2}}v^{{{\alpha}}{{\dot{\alpha}}}}z_{{\alpha}}^{\pm}\ ,\\[1.0pt] \nabla_{y^{{\dot{\alpha}}}_{\pm}}=\partial_{y^{{\dot{\alpha}}}_{\pm}}+{\textstyle\frac{1}{2}}v_{{{\dot{\alpha}}}{{\alpha}}}\bar{y}^{{\alpha}}_{\pm}\ ,\ \nabla_{\bar{y}^{{\alpha}}_{\pm}}=\partial_{\bar{y}^{{\dot{\alpha}}}_{\pm}}-{\textstyle\frac{1}{2}}v_{{{\dot{\alpha}}}{{\alpha}}}y^{{\dot{\alpha}}}_{\pm}\ ,\end{array}\end{split}$ (5.4)
@@ 5.3. Klein-Gordon, Dirac and Proca equations
The Lorentz conditions in (5.28) follow from the requrement of symmetry with respect to indices $i_{1},i_{2}$ in (4.58) and remove the extra component from $B_{\mu}^{\pm}$. This is the component 1 in the decomposition $({\textstyle\frac{1}{2}},{\textstyle\frac{1}{2}})={\bf 3}\oplus{\bf 1}$, when the group $\mathrm{SL}(2,\mathbb{C})$ is reduced to SU(2). The fields $B_{c}^{\pm}:=(-\mathrm{i}{{\gamma}}^{2})(B^{\pm})^{*}\mathrm{i}{{\gamma}}^{2}=(B_{\mu}^{\pm})^{*}{{\gamma}}^{\mu}$ are charge conjugate for the complex fields (5.29). The Proca field is called neutral (i.e. the particle coincides with antiparticle), if $B^{-}=B_{c}^{+}$. Then the field $B_{\mu}=B_{\mu}^{+}+B_{\mu}^{-}$ is real.
@@ 6.2. Photons in position representation
The first two equations in (6.11) are Maxwell’s equations (6.5) in the Lorentz gauge both for the field $A_{\mu}^{R}$ and the field $A_{\mu}^{L}$. The last equations in (6.11) mean that the complex gauge field $F_{\mu\nu}^{R}=\partial_{\mu}A_{\nu}^{R}-\partial_{\nu}A_{\mu}^{R}$ is self-dual, and the field $F_{\mu\nu}^{L}=\partial_{\mu}A_{\nu}^{L}-\partial_{\nu}A_{\mu}^{L}$ is anti-self-dual. Therefore, the solutions of equation (6.10) define complex self-dual and anti-self-dual solutions of Maxwell’s equations. Thus, the complex vector fields $A_{\mu}^{L}$ and $A_{\mu}^{R}$ introduced by us are photons in the Penrose sense, but defined in terms of gauge potentials in accordance with QFT. Note that in the case of massless fields we should not speak about spin, but about helicity, which in the case under consideration coincides with chirality. The self-dual field $F^{R}$ has helicity $h=1$, and the anti-self-dual field $F^{L}$ has helicity h=1h=-1.
@@ 7. Summary
$\omega_{\mathbb{R}^{3}}=-\mathrm{i}\frac{s}{2R^{3}}{\varepsilon}_{abc}S^{a}\mathrm{d}S^{b}\wedge\mathrm{d}S^{c}\ {\ \mathrm{for}\ }\ R^{2}:=\delta_{ab}S^{a}S^{b}\ ,$ (7.1)

### v2

@@ 1. Introduction
In fact, already at the level of classical mechanics one should understand what the spin of a particle is and what an antiparticle is. The situation with spin is as follows. The relativistic phase space $T^{*}\mathbb{R}^{1,3}$ of a massive particle with fixed parity extends to a phase space $T^{*}\mathbb{R}^{1,3}\times\mathbb{C}^{2}$, where $\mathbb{C}^{2}$ is the fundamental representation of the group $\mathrm{SL}(2,\mathbb{C})$ double covering the proper ortochronous Lorentz group $\mathrm{SO}^{+}(1,3)$ [1]. Next, a formal transition is made to the orbit $T^{*}H_{+}^{3}\times{\mathbb{C}P}^{1}$ of the Poincare group, over this orbit a complex line bundle $L_{\mathbb{C}}^{+}$ is introduced and the polarized sections of this bundle define fields with spin $s=0,{\textstyle\frac{1}{2}},1,$... [1]. Here $H_{+}^{3}$ is one sheet of a two-sheeted hyperboloid $\eta^{\mu\nu}p_{\mu}p_{\nu}=m^{2}$ in the momentum space, $p_{\mu}\in\mathbb{R}^{1,3}$, $\mu,\nu,...=0,...,3$, $(\eta_{\mu\nu})={\mathrm{diag}}(1,-1,-1,-1)$. If the parity is not fixed, then the phase space should be extended to the space $T^{*}\mathbb{R}^{1,3}\times\mathbb{C}^{2}_{L}\times\mathbb{C}_{R}^{2}$, where $\mathbb{C}_{L}^{2}$ and $\mathbb{C}_{R}^{2}$ are the representation spaces of the Lorentz group of type $({\textstyle\frac{1}{2}},0)$ and $(0,{\textstyle\frac{1}{2}})$. This allows us to obtain a description of fields in arbitrary representations $(s,j)$ of the Lorentz group, and not only representations of type $(s,0)$ and $(0,s)$ obtained for fixed parity. In the non-relativistic case, the phase space $\mathbb{C}^{2}_{L}\times\mathbb{C}_{R}^{2}$ is reduced to the fundamental representation $\mathbb{C}^{2}$ of the group SU(2) covering the rotation group SO(3), and the coset space ${\mathbb{C}P}^{1}=\mathrm{SU}(2)/\mathrm{U}(1)\subset\mathbb{C}^{2}$ of this group is used to introduce the space $\mathbb{C}^{2s+1}$ of quantum spin. The space of quantum spin in both the relativistic and non-relativistic cases is defined as the space $\mathbb{C}^{2s+1}$ of sections of the bundle $\mathcal{O}(2s)$ over the space ${\mathbb{C}P}^{1}$.
@@ 2.1. Dynamical spin variables
$N_{int}^{L}:=\frac{p^{{{\alpha}}\dot{{\alpha}}}}{m}z_{{\alpha}}\bar{z}_{\dot{{\alpha}}}=2s\ ,$ (2.6)
@@ 2.3. Spinning particles
$V_{S^{a}}=\frac{1}{2\mathrm{i}}\,\sigma^{{\alpha}}_{a\beta}\left(z_{{\alpha}}\frac{\partial}{\partial z_{\beta}}+\hat{z}^{{\alpha}}\frac{\partial}{\partial\hat{z}_{\beta}}\right)\quad\Rightarrow\quad[V_{S^{a}},V_{S^{b}}]={\varepsilon}^{c}_{ab}V_{S^{c}}\ ,$ (2.37)
@@ 2.5. Higher spins
$\mathrm{d}s^{2}_{\mathbb{C}^{2}/\mathbb{Z}_{n}}=\mathrm{d}\rho^{2}+\rho^{2}\mathrm{d}s^{2}_{S^{3}/\mathbb{Z}_{n}}=\mathrm{d}\rho^{2}+\frac{\rho^{2}}{n^{2}}(\mathrm{d}{{\varphi}}-\mathrm{i}a_{-n})^{2}+\frac{\rho^{2}\mathrm{d}z\mathrm{d}{\bar{z}}}{(1+z{\bar{z}})^{2}}\ .$ (2.53)
@@
$\mathrm{d}s^{2}_{\mathbb{C}/\mathbb{Z}_{m}}=\mathrm{d}z\mathrm{d}{\bar{z}}|_{\mathbb{C}/\mathbb{Z}_{m}}=\frac{1}{m^{2}}(\psi_{m}^{*}\psi_{m})^{\frac{1-m}{m}}\mathrm{d}\psi_{m}\mathrm{d}\psi_{m}^{*}=\mathrm{d}\rho^{2}+\frac{\rho^{2}}{m^{2}}\mathrm{d}{{\varphi}}^{2}\ ,$ (2.56)
@@ 3.3. Schrödinger equations
$\nabla_{x^{a}}=\frac{\partial}{\partial x^{a}}-p_{a}{J_{\sf{v}}}\ ,\quad\nabla_{p_{a}}=\frac{\partial}{\partial p_{a}}\quad\Rightarrow\quad\hat{x}^{a}={J_{\sf{v}}}\nabla_{p_{a}}{\quad\mathrm{and}\quad}\hat{p}_{a}=\mathrm{i}\nabla_{x^{a}}\ .$ (3.24)
@@
$\nabla_{x^{a}}^{\pm}=\mp\mathrm{i}p_{a}\ ,\quad\nabla_{p_{a}}^{\pm}=\frac{\partial}{\partial p_{a}}\quad\Rightarrow\quad\hat{x}^{a}_{\pm}=\pm\mathrm{i}\frac{\partial}{\partial p_{a}}{\quad\mathrm{and}\quad}\hat{p}_{a}^{\pm}=\pm p_{a}\ ,$ (3.25)
@@ 3.4. Wave function collapse
Thus, the probability of detecting certain parameters $(s,s_{3})$ in an experiment arises in the case where we do not know completely the initial conditions of the particle’s motion. This is equivalent to how probabilities arise when throwing dice.
@@ 3.5. From Schrödinger to Pauli equation
$\langle\Psi_{\pm},\Psi_{\pm}\rangle=\Phi_{\pm}^{\dagger}\Phi_{\pm}=\delta_{{{\alpha}}{{\dot{\alpha}}}}\psi_{\pm}^{{\alpha}}\bar{\psi}_{\pm}^{{\dot{\alpha}}}{\quad\mathrm{and}\quad}\rho_{\pm}=\pm\Phi_{\pm}^{\dagger}\Phi_{\pm}$ (3.57)
@@ 4. Relativistic Hamiltonian mechanics
4.1. Klein-Gordon particles
@@ 4.4. Dirac and Proca particles
$g_{D}=-\mathrm{i}\,\mathrm{d}Z^{\pm}J\mathrm{d}\bar{Z}^{\pm}=\mathrm{d}Z^{+}\mathrm{d}\bar{Z}^{+}=-\mathrm{d}Z^{-}\mathrm{d}\bar{Z}^{-}=\mathrm{d}z_{{\alpha}}\mathrm{d}\bar{y}^{{\alpha}}+\mathrm{d}y^{{\dot{\alpha}}}\mathrm{d}{\bar{z}}_{{\dot{\alpha}}}\ ,$ (4.60)
@@ 5.2. Relativistic analogue of Schrödinger equation
$\begin{split}L_{\mathbb{C}}^{\pm}&\ \to\ T^{*}\mathbb{R}^{1,3}:\quad\nabla_{x^{\mu}}=\frac{\partial}{\partial x_{\mu}}\pm\mathrm{i}p_{\mu}\ ,\quad\nabla_{p_{\mu}}=\frac{\partial}{\partial p_{\mu}}\ ,\\[2.0pt] L_{\mathbb{C}}^{\pm}&\ \to\ \mathbb{C}_{\pm}^{4}:\quad\begin{array}[]{l}\nabla_{z_{{\alpha}}^{\pm}}=\partial_{z_{{\alpha}}^{\pm}}+{\textstyle\frac{1}{2}}v^{{{\alpha}}{{\dot{\alpha}}}}{\bar{z}}_{{\dot{\alpha}}}^{\pm}\ ,\ \nabla_{{\bar{z}}_{{\dot{\alpha}}}^{\pm}}=\partial_{{\bar{z}}_{{\dot{\alpha}}}^{\pm}}-{\textstyle\frac{1}{2}}v^{{{\alpha}}{{\dot{\alpha}}}}z_{{\alpha}}^{\pm}\ ,\\[1.0pt] \nabla_{y^{{\dot{\alpha}}}_{\pm}}=\partial_{y^{{\dot{\alpha}}}_{\pm}}+{\textstyle\frac{1}{2}}v_{{{\dot{\alpha}}}{{\alpha}}}\bar{y}^{{\alpha}}_{\pm}\ ,\ \nabla_{\bar{y}^{{\alpha}}_{\pm}}=\partial_{\bar{y}^{{\dot{\alpha}}}_{\pm}}-{\textstyle\frac{1}{2}}v_{{{\dot{\alpha}}}{{\alpha}}}y^{{\dot{\alpha}}}_{\pm}\ ,\end{array}\end{split}$ (5.4)
@@ 5.3. Klein-Gordon, Dirac and Proca equations
The Lorenz conditions in (5.28) follow from the requirement of symmetry with respect to indices $i_{1},i_{2}$ in (4.58) and remove the extra component from $B_{\mu}^{\pm}$. This is the component 1 in the decomposition $({\textstyle\frac{1}{2}},{\textstyle\frac{1}{2}})={\bf 3}\oplus{\bf 1}$, when the group $\mathrm{SL}(2,\mathbb{C})$ is reduced to SU(2). The fields $B_{c}^{\pm}:=(-\mathrm{i}{{\gamma}}^{2})(B^{\pm})^{*}\mathrm{i}{{\gamma}}^{2}=(B_{\mu}^{\pm})^{*}{{\gamma}}^{\mu}$ are charge conjugate for the complex fields (5.29). The Proca field is called neutral (i.e. the particle coincides with antiparticle), if $B^{-}=B_{c}^{+}$. Then the field $B_{\mu}=B_{\mu}^{+}+B_{\mu}^{-}$ is real.
@@ 6.2. Photons in position representation
The first two equations in (6.11) are Maxwell’s equations (6.5) in the Lorenz gauge both for the field $A_{\mu}^{R}$ and the field $A_{\mu}^{L}$. The last equations in (6.11) mean that the complex gauge field $F_{\mu\nu}^{R}=\partial_{\mu}A_{\nu}^{R}-\partial_{\nu}A_{\mu}^{R}$ is self-dual, and the field $F_{\mu\nu}^{L}=\partial_{\mu}A_{\nu}^{L}-\partial_{\nu}A_{\mu}^{L}$ is anti-self-dual. Therefore, the solutions of equation (6.10) define complex self-dual and anti-self-dual solutions of Maxwell’s equations. Thus, the complex vector fields $A_{\mu}^{L}$ and $A_{\mu}^{R}$ introduced by us are photons in the Penrose sense, but defined in terms of gauge potentials in accordance with QFT. Note that in the case of massless fields we should not speak about spin, but about helicity, which in the case under consideration coincides with chirality. The self-dual field $F^{R}$ has helicity $h=1$, and the anti-self-dual field $F^{L}$ has helicity h=1h=-1.
@@ 7. Summary
$\omega_{\mathbb{R}^{3}}=-\frac{s}{R^{3}}\,{\varepsilon}_{abc}\,S^{a}\mathrm{d}S^{b}\wedge\mathrm{d}S^{c}\ {\ \mathrm{for}\ }\ R^{2}:=\delta_{ab}S^{a}S^{b}\ ,$ (7.1)